### v1

@@ Abstract
We consider a mean-field optimal control problem with general dynamics including common noise and jumps and show that its minimizers are Nash equilibria of an associated mean-field game of controls. These types of games are necessarily potential, and the Nash equilibria obtained as the minimizers of the control problems are naturally related to the McKean-Vlasov equations of Langevin type. We provide several examples to motivate the general theory including the Cucker-Smale Flocking and Kuramoto mean-field games. The invariance property of the value function of these control problems, utilized in our proof, is also proved.
@@ 1 Introduction
Mean-field games (MFGs) were independently introduced by Lasry & Lions [46, 47, 48] and Huang, Caines, & Malhamé [35, 36, 37, 38] to approximate large population games in which the interaction appears through the empirical distribution of all agents. Since their introduction, they have been widely applied in contexts ranging from heterogeneous agents models in economics to models of jet-lag recovery in neuroscience.
@@
We study a special class of MFGs that arise as first-order conditions of a mean-field control problem (MFC). This connection has first been observed in the seminal work [48] by Lasry & Lions and has been subsequently well-studied under the name potential mean-field games as we review below. Our primary aim is to explore the natural links between the potential MFG and MFC problems as well as the links with a class of Wasserstein gradient flows. The latter connection can be leveraged to construct compelling mean-field games that have similar qualitative behavior as the gradient flows, and in Section 6 we discuss two such examples: the MFG flocking model of [54, 55] and the Kuramoto MFG of [14, 60, 61].
Specifically, let ${\mathscr{P}}_{2}(\mathbb{R}^{n})$ be the Wasserstein space of probability measures on $\mathbb{R}^{n}$ with finite second moments. For a sufficiently differentiable function $F:{\mathscr{P}}_{2}(\mathbb{R}^{n})\to\mathbb{R}$, consider the following McKean-Vlasov stochastic differential equation:
$\mathrm{d}X_{t}=-\nabla_{x}\delta_{\nu}F({\mathcal{L}}(X_{t}))(X_{t})\,\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (1.1)
where $\delta_{\mu}F(\mu)(x)$ is the linear derivative of $F$ at $\mu\in{\mathscr{P}}_{2}(\mathbb{R}^{n})$, ${\mathcal{L}}(X_{t})$ denotes the law of $X_{t}$, and $W$ is a standard Brownian motion in $\mathbb{R}^{n}$. In analogy to finite-dimensional Langevin dynamical systems, [34] refers to (1.1) as the mean-field Langevin dynamics.
Recently, these equations have received considerable attention due to their connection to noisy gradient descent algorithms in deep neural networks [19, 34, 52, 53]. Additionally, we argue in Section 6 that one can construct potential MFC problems with similar qualitative behavior. In particular, the optimally controlled state dynamics are in the above form (1.1).
While most existing studies on potential games, such as [4], start from a MFG and establish structural conditions on the given functions under which equilibria can be obtained from a MFC problem, we start from a general MFC problem and construct an associated potential MFG. In summary, we propose the following relations,
MF-Langevin Dynamics MFC Problem Potential MFG
Additionally, by working with the probabilistic formulation of both the control and game problems, we obtain general results with non-Markovian controlled dynamics under jumps and common noise. Our approach extends to cost functions that depend on the joint distribution of the state and control and which are non-separable in the control and measure variable. This further allows us to derive mean-field games of controls.
MFC, MFG, and Main Results.
@@
On a filtered probability space $(\Omega,{\mathscr{F}},\mathbb{F},\mathbb{P})$ satisfying the usual conditions, let $W^{0}$ and $W$ be $\mathbb{F}$-Brownian motions, and $N$ and $N^{0}$ be stationary $\mathbb{F}$-Poisson random measures on $[0,\infty)\times(\mathbb{R}^{k}\setminus\{0\})$. In our setting, $W$ and $N$ represent the idiosyncratic noise, while $W^{0}$ and $N^{0}$ model the common noise. Let $T>0$ be the time horizon. For a given initial condition $X_{0}$, and coefficients $(b,\sigma,\sigma^{0},\lambda,\lambda^{0})$ satisfying Assumption 2.3 below, the controlled state process $X^{\alpha}\in\mathbb{R}^{n}$ is the unique strong solution to the following stochastic differential equation,
$X^{\alpha}_{t}=$ $X_{0}+\int_{0}^{t}b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\left(\sigma(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W_{s}+\sigma^{0}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W^{0}_{s}\right)$ (1.2)
@@
$+\int_{0}^{t}\int_{\mathbb{R}^{k}\setminus\{0\}}\left(\lambda(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{N}(\mathrm{d}s,\mathrm{d}\zeta)+\lambda^{0}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\ \widetilde{N}^{0}(\mathrm{d}s,\mathrm{d}\zeta)\right),$
@@
where $\alpha_{\cdot}$ is any $\mathbb{F}$-predictable, $\mathrm{d}\mathbb{P}\otimes\mathrm{d}t$-square integrable control process taking values in a Borel subset $U$ of some Euclidean space. Let ${\mathscr{A}}$ denote the set of all such control processes. Notice that the above dynamics do not contain any mean field interactions, and this is our main structural assumption.
@@
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)\mapsto\mathbb{R},\quad G:{\mathscr{D}}_{n}\times{\mathscr{P}}({\mathscr{D}}_{n})\mapsto\mathbb{R},$
where ${\mathscr{D}}_{n}$ is the Skorokhod space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{n}$, and for any measurable space $S$, ${\mathscr{P}}(S)$ is the set of probability measures on $S$. We set
@@
where ${\mathcal{L}}^{\alpha}_{t}:={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,W^{0},N^{0})$ is the regular conditional distribution of $(X^{\alpha}_{\cdot\land t},\alpha_{t})$ given the common noise $W^{0},N^{0}$. Then, the mean-field control (MFC) problem consists of minimizing $J(X_{0},\alpha)$ over all $\alpha\in{\mathscr{A}}$.
We next introduce the associated mean-field game of controls. Let $\Pi:{\mathscr{D}}_{n}\times U\mapsto{\mathscr{D}}_{n}$ be the projection on the first coordinate, and for $(t,x,a,\nu)\in[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)$, let $\mu=\nu\circ\Pi^{-1}$ and set
@@
Given an adapted stochastic process $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ with values in ${\mathscr{P}}({\mathscr{D}}_{n}\times U)$, we set
$J_{g}(X_{0},\alpha,\nu^{*}):=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t+g(X^{\alpha}_{\cdot\land T},\nu^{*}_{T})\right].$
@@
The precise formulation is given in Theorem 3.6 below. Additionally, in Section 3 and subsection 4.3 below, we discuss specific examples and structures, and show that this theorem covers all previous results. In particular, it is clear that the results of this paper remain true in the absence of the common noise by simply setting $\sigma^{0}=\lambda^{0}=0$ and in the absence of jumps by setting $\lambda=\lambda^{0}=0$. Moreover, as discussed in subsection 4.2, in the classical separable models, we obtain a standard MFG depending only on the distribution of the state process and not the control.
@@
One technical tool we utilize is the invariance principle Theorem 7.2. It states that the value function of MFC and the Nash equilibria of MFG do not depend on the underlying probabilistic structure. This property is needed in our approach since in our proof we expand the initially given probabilistic structure by adding independent Bernoulli random variables. For diffusions it is established in [20, 25], and we extend it to our setting.
Context and Earlier Results.
@@
Succeeding the introduction of potential MFGs in [48, Section 2.6], they have been studied in [4, 6, 7, 10, 56] in both the first-order (i.e. without Brownian noise) and the second-order cases. The majority of the existing literature on potential mean-field games starts from a characterization of the Nash equilibria as solutions of a forward-backward system of partial differential equations. These works interpret the PDE system as an optimality condition for two problems in duality by using an elegant application of the Fenchel-Rockafellar Duality Theorem or the von Neumann Minimax Principle, as illustrated in [56]. For the first-order potential MFGs we refer readers to the recent paper [30] which provides an excellent introduction to these techniques and also a historical context. A similar analysis in the context of optimal transport can be found in [9] and for (first-order) variational MFGs in [1, Section 2]. Extensions of these results include infinite-horizon cost criteria [51] and problems with state constraints [22]. The structure of potential MFGs has been leveraged to study weak solutions of the master equation [16] and to prove the convergence of learning procedures to solutions of the MFG system [4, 8]. We also refer to [32] for the study of stochastic differential potential games with finitely many players.
@@
On finite state spaces, direct arguments such as the Pontryagin principle imply that potential MFG equilibria can be viewed as first-order conditions of an associated control problem as initiated in [29, 31]. For MFGs with finite state space, the long-time behavior is investigated in [27]; the convergence problem of the $N$-player game in a two-state model without uniqueness is examined in [17]. For finite state problems with no uniqueness, [15] proposes a selection principle of MFG equilibria as the minimizers of the associated control problem when uniqueness fails. In the continuous state case, an approach to potential MFGs that is not based on convex duality is presented in the classical book [11]. Chapters 6.2.5 and 6.7.2 of [11] identify the Pontryagin system of the MFC problem with the FBSDE system describing MFG Nash equilibria in the probabilistic setting.
@@
We work in a general stochastic formulation of the control problem using direct methods, rather than the FBSDEs. Our results are in considerable generality, extending to $(i)$ the environments with common noise, $(ii)$ the models with jumps, $(iii)$ the cost functionals $F,G$ that depend on the control and measure components in a non-separable way, and $(iv)$ coefficients that depend on the entire path of the state variable. The non-separability of the cost functions leads to a mean-field game of controls, where the interaction between agents is through both the state and control variables. Further discussion is given in subsection 4.3.
@@
The paper is organized as follows: After reviewing the notation and the setting, in Section 3 we define the MFC problem and the MFG of controls, and then state the main result. We discuss the result and several special cases in Section 4, and the proof of our main result Theorem 3.6 is given in Section 5. We propose a connection between the Langevin dynamics, MFC, and MFG, and illustrate it through two examples in Section 6. Section 7 proves the invariance principle. Appendix A provides a summary of standard results and notations on point processes, and Appendix B discusses the discounted infinite horizon setting.
@@ 2 Notation and Setting
For a topological space $S$, ${\mathscr{B}}(S)$ is the set of all Borel sets. The dimension of the state process is denoted by $n$, $d$ refers to the dimension of the idiosyncratic Brownian motion, and $\ell$ is the dimension of the common Brownian noise. $T>0$ is the horizon. ${\mathscr{D}}_{k}(0,T)$ is the Polish space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{k}$ endowed it with the Skorokhod $J_{1}$ topology. When the domain is clear, we write ${\mathscr{D}}_{k}$. On ${\mathscr{D}}_{k}$, we sometimes utilize the supremum norm $\|x\|_{T}:=\sup_{t\in[0,T]}|x(t)|$. The control set $U$ is a non-empty Borel subset of a Euclidean space. On any Polish space $S$, ${\mathscr{P}}(S)$ is the set of probability measures, and ${\mathscr{P}}_{2}(S)$ is the Wasserstein space of measures with finite second moments. A function $\phi:{\mathscr{P}}_{2}(S)\to\mathbb{R}$ is called (linearly) differentiable, if there exists a measurable function $\delta_{\eta}\phi:{\mathscr{P}}_{2}(S)\times S\to\mathbb{R}$ which has at most of quadratic growth in $x\in S$ and satisfies,
@@
For a $\sigma$-algebra ${\mathscr{G}}$ and an $S$-valued random variable $Y$, the random measure ${\mathcal{L}}(Y\,|\,{\mathscr{G}})\in{\mathscr{P}}(S)$ is the regular condition distribution of $Y$ given ${\mathscr{G}}$. We write ${\mathcal{L}}^{\mathbb{P}}={\mathcal{L}}$, when the dependence on the underlying probability measure $\mathbb{P}$ is important. When ${\mathscr{G}}$ is trivial, ${\mathcal{L}}(Y)$ is the law of $Y$. Finally, ${\mathscr{L}}^{2}({\mathscr{G}})$ denotes the set of all square-integrable $n$-dimensional random variables which are ${\mathscr{G}}$-measurable.

### v2

@@ Abstract
We study a mean field optimal control problem with general non-Markovian dynamics, including both common noise and jumps. We show that its minimizers are Nash equilibria of an associated mean field game of controls. These types of games are necessarily potential, and the Nash equilibria derived as the minimizers of the control problem are closely connected to McKean-Vlasov equations of Langevin type. To illustrate the general theory, we present several examples, including a mean field game of controls with interactions through a price variable, and mean field Cucker-Smale Flocking and Kuramoto models. We also establish the invariance property of the value function, a key ingredient used in our proofs.
@@ 1 Introduction
Mean field games (MFGs) were independently introduced by Lasry & Lions [49, 50, 51] and Huang, Caines, & Malhamé [39, 40, 41, 42] as a framework to approximate large population games in which the interaction appears through the empirical distribution of agents. Since their introduction, they have been widely applied in contexts ranging from heterogeneous agent models in economics to models of jet-lag recovery in neuroscience.
@@
In this work, we focus on a particular class of MFGs that emerge as the first-order optimality condition of a mean field control problem (MFC). This connection was first identified in the seminal paper [51] by Lasry & Lions and has since been extensively studied under the name potential mean field games as we review below. Our primary goal is to establish a general connection between potential MFG and MFC problems. In addition, we identify a close link to a class of Wasserstein gradient flows. The latter connection can be leveraged to construct compelling mean field games that have similar qualitative behavior as the gradient flows. In Section 6, we discuss two such examples: the MFG flocking model of [57, 58] and the Kuramoto MFG of [16, 63, 64].
While most existing studies on potential games, e.g. [6], start from a MFG and establish structural conditions on the cost functions under which equilibria can be obtained from a MFC problem, we start from a general MFC problem and construct an associated potential MFG. By adopting a probabilistic formulation for both the control and game problems, we obtain general results that accommodate non-Markovian dynamics under jumps and common noise. In the case of separable cost structures, our framework recovers existing results, while in the non-separable case, it leads to mean field games of controls.
1.1 MFC, MFG and Main Results
@@
On a filtered probability space $(\Omega,{\mathscr{F}},\mathbb{F},\mathbb{P})$ satisfying the usual conditions, let $W^{0}$ and $W$ be $\mathbb{F}$-Brownian motions, and $N$ and $N^{0}$ be stationary $\mathbb{F}$-Poisson random measures on $[0,\infty)\times(\mathbb{R}^{k}\setminus\{0\})$. In our setting, $W$ and $N$ represent the idiosyncratic noise, while $W^{0}$ and $N^{0}$ model the common noise. We write $\widetilde{N}$ and $\widetilde{N}^{0}$ for the corresponding compensated random measures. Let $T>0$ be the time horizon. For a given initial condition $X_{0}$, and coefficients $(b,\sigma,\sigma^{0},\lambda,\lambda^{0})$ satisfying Assumption 2.3 below, the controlled state process $X^{\alpha}\in\mathbb{R}^{n}$ is the unique strong solution to the following stochastic differential equation,
@@
$\begin{split}X^{\alpha}_{t}=&X_{0}+\int_{0}^{t}b(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}s+\int_{0}^{t}\left(\sigma(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W_{s}+\sigma^{0}(s,X^{\alpha}_{\cdot\land s},\alpha_{s})\,\mathrm{d}W^{0}_{s}\right)\\ &+\int_{0}^{t}\int_{\mathbb{R}^{k}\setminus\{0\}}\left(\lambda(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\widetilde{N}(\mathrm{d}s,\mathrm{d}\zeta)+\lambda^{0}(s,X^{\alpha}_{\cdot\land s-},\alpha_{s},\zeta)\,\ \widetilde{N}^{0}(\mathrm{d}s,\mathrm{d}\zeta)\right),\end{split}$ (1.1)
@@
where $\alpha_{\cdot}$ is any $\mathbb{F}$-predictable, $\mathrm{d}\mathbb{P}\otimes\mathrm{d}t$-square integrable control process taking values in a Borel subset $U$ of some Euclidean space. Let ${\mathscr{A}}$ denote the set of all such control processes. Notice that the above dynamics do not contain any mean field interactions. We refer to Section 3.4 for a class of control problems in which some of the dynamics are allowed to be of mean field type.
@@
$F:[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)\to\mathbb{R},\quad G:{\mathscr{D}}_{n}\times{\mathscr{P}}({\mathscr{D}}_{n})\to\mathbb{R},$
where ${\mathscr{D}}_{n}$ denotes the Skorokhod space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{n}$, and for any measurable space $S$, ${\mathscr{P}}(S)$ is the set of probability measures on $S$. We set [^4] We note that the terminal cost $G$ depends on ${\mathcal{L}}^{\alpha}_{T}$ only through its first marginal.
@@
where ${\mathcal{L}}^{\alpha}_{t}:={\mathcal{L}}((X^{\alpha}_{\cdot\land t},\alpha_{t})\,|\,W^{0},N^{0})$ is the regular conditional distribution of $(X^{\alpha}_{\cdot\land t},\alpha_{t})$ given the common noise $(W^{0},N^{0})$. Then, the mean field control (MFC) problem consists of minimizing $J(X_{0},\alpha)$ over all $\alpha\in{\mathscr{A}}$.
We next introduce the associated mean field game of controls. Let $\Pi:{\mathscr{D}}_{n}\times U\to{\mathscr{D}}_{n}$ be the projection on the first coordinate, and for $(t,x,a,\nu)\in[0,T]\times{\mathscr{D}}_{n}\times U\times{\mathscr{P}}({\mathscr{D}}_{n}\times U)$, let $\mu=\Pi_{\sharp}\nu$ and set
@@
Given an adapted stochastic process $\nu^{*}=(\nu^{*}_{t})_{t\in[0,T]}$ with values in ${\mathscr{P}}({\mathscr{D}}_{n}\times U)$, we define the game cost of a typical agent by,
$J_{g}(X_{0},\alpha,\nu^{*}):=\mathbb{E}\left[\int_{0}^{T}f(t,X^{\alpha}_{\cdot\wedge t},\alpha_{t},\nu^{*}_{t})\,\mathrm{d}t+g(X^{\alpha}_{\cdot\land T},\nu^{*}_{T})\right].$ (1.4)
@@
The precise formulation is provided in Theorem 3.6 below. Additionally, in Sections 3 and 5.4 below, we discuss specific structures which show that this theorem covers previously known results in the literature where dynamics are independent of the law. An extension to law-dependent dynamics is treated in Subsection 3.4. Furthermore, it is clear that the results of this paper remain true in the absence of the common noise by simply setting $\sigma^{0}=\lambda^{0}=0$ and in the absence of jumps by setting $\lambda=\lambda^{0}=0$. As discussed in Subsection 5.2, in the classical separable models, we obtain a standard MFG depending only on the distribution of the state process and not the control.
@@
A key technical tool in our analysis is the invariance principle presented in Theorem 7.3. This result states that both the value function of the MFC problem and the Nash equilibria of the MFG are independent of the underlying probabilistic structure. This property is essential to our approach, as our proof requires extending the original probabilistic basis by adding independent Bernoulli random variables. The invariance principle has been established for diffusions in [23, 28], and we extend it to our setting.
1.2 Context and Earlier Results
@@
Following the introduction of potential mean field games (MFGs) in [51, Section 2.6], they have been developed in [6, 8, 9, 12, 59, 34, 17] in both the first-order (i.e., without Brownian noise) and second-order settings. The majority of the existing literature on potential mean field games starts from a characterization of the Nash equilibria as solutions of a forward-backward system of partial differential equations. These works interpret the PDE system as an optimality condition by using an elegant application of the Fenchel-Rockafellar Duality Theorem or the von Neumann Minimax Principle, as illustrated in [59].
For a comprehensive introduction to these techniques — especially in the first-order case — we refer the reader to the recent work [32], which also provides historical context. Related analyses appear in the context of optimal transport in [11] and for (first-order) variational MFGs in [1, Section 2]. Extensions of these results include infinite-horizon cost criteria [54] and problems with state constraints [25]. The structure of potential MFGs has been leveraged to study weak solutions of the master equation [19] and to prove the convergence of learning procedures to solutions of the MFG system [6, 10]. Applications of potential MFGs include games for market competition [33, 5]. We also refer to [36] for the study of stochastic differential potential games with finitely many players.
@@
On finite state spaces, direct arguments such as the Pontryagin principle imply that potential MFG equilibria can be viewed as first-order conditions of an associated control problem as initiated in [31, 35]. For finite state problems with no uniqueness, [18] proposes a selection principle of MFG equilibria as the minimizers of the associated control problem when uniqueness fails. In the continuous state case, an approach to potential MFGs that is not based on convex duality is presented in the classical book [13]. Chapters 6.2.5 and 6.7.2 of [13] identify the Pontryagin system of the MFC problem with the FBSDE system describing MFG Nash equilibria in the probabilistic setting. The forthcoming paper [17] discusses potential mean field games in the weak formulation.
@@
We work in a general stochastic formulation of the control problem using direct methods, rather than the FBSDEs. Our results are in considerable generality, extending to problems featuring $(i)$ common noise, $(ii)$ jumps, $(iii)$ non-separable cost structures and $(iv)$ coefficients that depend on the entire path of the state variable. The non-separability of the cost functions leads to a mean field game of controls, where the interaction between agents is through both the state and control variables. Potential MFGs of controls have been studied in [34, 5] in their analytic formulation. We provide further discussion of the deterministic and probabilistic formulations in Subsection 5.4.
@@
The paper is organized as follows. In Section 3, after reviewing notation, we define the mean field control (MFC) problem and the mean field game (MFG) of controls, and we present our main result. We then discuss the result and several special cases in Section 5, and the proof of our main result Theorem 3.6 is given in Section 4. Examples of potential mean field games of controls and the connection between the Langevin dynamics, MFC, and MFG are presented in Section 6. Section 7 proves the invariance principle. Appendix A provides a summary of standard results and notations on point processes, and Appendix B discusses the discounted infinite horizon setting.
@@ 2 Notation and Setting
For a topological space $S$, ${\mathscr{B}}(S)$ is the set of all Borel sets. The dimension of the state process is denoted by $n$, $d$ refers to the dimension of the idiosyncratic Brownian motion, and $\ell$ is the dimension of the common Brownian noise. $T>0$ is the horizon. ${\mathscr{D}}_{k}(0,T)$ is the Polish space of càdlàg functions from $[0,T]$ to $\mathbb{R}^{k}$ endowed it with the Skorokhod $J_{1}$ topology. When the domain is clear, we write ${\mathscr{D}}_{k}$. On ${\mathscr{D}}_{k}$, we sometimes utilize the supremum norm $\|x\|_{T}:=\sup_{t\in[0,T]}|x(t)|$. The control set $U$ is a non-empty Borel subset of a Euclidean space. On any Polish space $S$, ${\mathscr{P}}(S)$ is the set of probability measures, and ${\mathscr{P}}_{2}(S)$ is the Wasserstein space of measures with finite second moments. We will frequently consider measures on the product space ${\mathscr{D}}_{k}\times U$ and will generically denote by $\Pi(x,a)=x$, $(x,a)\in{\mathscr{D}}_{k}\times U$, the projection on the first coordinate. If $\nu\in{\mathscr{P}}({\mathscr{D}}_{k}\times U)$, then $\Pi_{\sharp}\nu$ refers to the push-forward of $\nu$ under $\Pi$. A function $\phi:{\mathscr{P}}_{2}(S)\to\mathbb{R}$ is called (linearly) differentiable, if there exists a measurable function $\delta_{\eta}\phi:{\mathscr{P}}_{2}(S)\times S\to\mathbb{R}$ which is of at most quadratic growth in $x\in S$ and satisfies,
@@
If $\phi:{\mathscr{P}}_{2}(S)\to\mathbb{R}^{k}$ is vector-valued, then we say that $\phi$ is differentiable if all its components are differentiable. In this case we set, $\delta_{\eta}\phi=(\delta_{\eta}\phi_{1},\ldots,\delta_{\eta}\phi_{k})$. For a $\sigma$-algebra ${\mathscr{G}}$ and an $S$-valued random variable $Y$, the random measure ${\mathcal{L}}(Y\,|\,{\mathscr{G}})\in{\mathscr{P}}(S)$ is the regular condition distribution of $Y$ given ${\mathscr{G}}$. We write ${\mathcal{L}}^{\mathbb{P}}={\mathcal{L}}$, when the dependence on the underlying probability measure $\mathbb{P}$ is important. When ${\mathscr{G}}$ is trivial, ${\mathcal{L}}(Y)$ is the law of $Y$. Finally, ${\mathscr{L}}^{2}({\mathscr{G}})$ denotes the set of all square-integrable $n$-dimensional random variables which are ${\mathscr{G}}$-measurable.